OpenStudy (anonymous):

\[\frac{ 2 }{ 7 } x= 14\]

OpenStudy (anonymous):

so \(x=14\times \frac{7}{2}\)

OpenStudy (skullpatrol):

Multiply both sides by 7/2.

OpenStudy (anonymous):

x=49

OpenStudy (skullpatrol):

\[\frac{ 7 }{ 2 } \times \frac{ 2 }{ 7 } = 1\]